# DivCon: Divide and Conquer for Progressive Text-to-Image Generation

### v1

@@ 4.1.1 Benchmarks
To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS[2] and NSR-1K[15]. The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships. The NSR-1K dataset consists of 762 prompts for numerical relationships and 283 prompts for spatial relationships, all labeled with names, counts, bounding boxes, and ground truth images. We also utilize NSR-1K for quality evaluation. Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations. This elevates HRS as a more sophisticated and challenging benchmark.
4.1.2 Evaluation Metrics
We evaluate the images produced by our methods in terms of both quality and fidelity.
@@ 4.1.2 Evaluation Metrics
(1) Quality To assess the quality of generated images, we use FID as the metric[18]. Lower FID represents higher similarity between generated images and ground truths images. We report FID in NSR-1K [15] as each prompt has a corresponding ground truth image for FID calculation.
@@
(2) Fidelity To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated. Particularly, YOLOv8 [20] is employed to obtain the bounding boxes of the objects in the generated images. In counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy [13, 16]. In spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills[11] and LayoutGPT [15]. Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship. Otherwise, they are denoted as a top-bottom relationship.
4.1.3 Baselines
Stable Diffusion [30], Attend-and-Excite[7], and Attention-Refocusing [26] are employed for comparison. Stable Diffusion and Attend-and-Excite directly generate images from the text prompt. Meanwhile, Attention-Refocusing employs layout as an intermedium for image generation.
Methods
Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc.
Spatial
Layout Image
Acc.
Acc.
Stable Diffusion 2.1[30]
- - - - 73.97 53.93 62.38 16.20
7.96
Attend-and-Excite[7]
- - - - 76.57 55.94 64.65 18.37
9.23
Attention-Refocusing[26]
88.62 91.91 90.44 77.05 77.43 61.78 68.72 25.60
89.57
48.29
DivCon (Ours)
94.17 89.53 91.80 80.73 79.13 63.53 70.48 29.62
91.87
53.87
Methods
Numerical Layout Pre. Rec. F1 Acc. Image Pre. Rec. F1 Acc.
Spatial
Layout Image
Acc.
Acc.
Stable Diffusion 2.1[30]
- - - - 78.94 75.42 77.41 33.99
22.97
Attend-and-Excite[7]
- - - - 79.21 78.62 78.91 39.51
29.61
Attention-Refocusing[26]
98.51 99.11 98.81 94.31 84.61 85.64 85.12 56.10
95.05
69.28
DivCon (Ours)
98.47 98.35 98.41 94.39 85.31 87.54 86.41 55.80
96.11
72.95
@@ Methods
Figure 6: Qualitative comparison of numerical reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing, and our DivCon.
Figure 7: Qualitative comparison of spatial reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing and our DivCon.
Table 3: FID results achieved by different methods in NSR-1K.
Methods
FID ($\downarrow$)
Stable Diffusion 2.1[30]
20.77
Attention-Refocusing[26]
21.05
DivCon (Ours)
20.89
@@ 4.2.1 Quantitative Results
We present the results on the HRS benchmark [2] in Tab. 1. As we can see, DivCon outperforms previous T2I models by over 10% in numerical accuracy and over 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance. In addition, DivCon surpasses Attention Refocusing [26] on most metrics. Specifically, our layout prediction improves numerical accuracy by 3% and spatial accuracy by 2%, respectively. For image generation, the F1 score and accuracy achieve 2% and 4% gains, while the spatial accuracy is boosted by up to 6%. By dividing the T2I task into simple subtasks, our DivCon can better understand the text prompt to achieve higher fidelity.
@@
The results on the NSR-1K benchmark are presented in Tab. 2. It can be observed that DivCon outperforms other T2I models by over 15% in numerical accuracy and over 45% in spatial accuracy. As compared to Attention Refocusing [26], our DivCon produces competitive results in terms of layout accuracy. Meanwhile, DivCon achieves approximately 2% superiority in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy.
Quantitative results in terms of image quality are shown in Tab. 3. As we can see, our DivCon produces compatible FID scores as compared to stable diffusion and outperforms attention-refocusing with notable margins. This demonstrates that our approach is able to preserve image quality while enhancing prompt fidelity.
4.2.2 Qualitative Results
@@
Fig. 6 and Fig. 7 illustrate the qualitative comparison of DivCon and baselines. In all cases, our DivCon can accurately predict the object quantities and the spatial relations from the text prompts.
@@
(1) Large Quantity of Objects. Given a prompt containing more than one object category (Fig. 6), Stable Diffusion [30] and Attend-and-Excite [7] often generate an insufficient quantity of objects, while Attention-Refocusing [26] sometimes omits one category of objects or generates an excessive number of objects. For instance, the prompt in Fig. 6(b) contains five spoons and four oranges. However, Stable Diffusion and Attend-and-Excite cannot synthesize sufficient spoons or oranges. Meanwhile, Attention-Refocusing generates an excess number of objects, with spoons appearing incomplete due to missing parts. As compared to these methods, our DivCon achieves superior numerical accuracy.
@@
(2) Complicated Spatial Relation. Given a prompt containing complex spatial relationships among more than two object categories (Fig. 7, previous T2I models struggle to generate correct spatial relations. Meanwhile, Attention-Refocusing tends to produce distortion or blending artifacts around the boundaries of objects. In Fig. 7 (b), Stable Diffusion and Attend-and-Excite miss certain object categories, while Attention-Refocusing merges the color of the banana with the person. In contrast, our DivCon inference accurate object categories and their spatial relationships, exhibiting superior performance in these challenging cases.
Figure 8: Comparison of layouts generated from Attention-Refocusing and our DivCon.
@@
Figure 10: Comparison of transmit one or two latent representations.
4.3 Ablation Study
In this sub-section, we conduct ablation experiments to evaluate the effectiveness of our divide-and-conquer approach in both layout prediction and image generation. Results are presented in Tab. 4 and Attention-Refocusing [26] is used as the baseline.
4.3.1 Divide-and-Conquer in Layout Prediction
To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step. As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 4% in both numerical and spatial accuracy. We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig. 8. Compared to Attention-Refocusing, layouts generated by our DivCon not only exhibit higher reasoning accuracy (Fig. 8 (a), (b)) but also are more organized with less overlap (Fig. 8 (c), (d)).
4.3.2 Divide-and-Conquer in Image Generation
@@
To study the effectiveness of our divide-and-conquer approach in the image generation stage, we develop a variant to synthesize the image in a single forward pass. It can be observed from Tab. 4 that the divide-and-conquer approach facilitates our method to notable performance gains in terms of all metrics. We further compare the results produced by our method using prompts containing multiple objects in Fig. 9. As we can see, the divide-and-conquer approach facilitates our DivCon to better reconstruct the objects, especially for small-sized ones, while avoiding generating objects outside the designated bounding boxes. For example, in Fig. 9(a), DivCon accurately reconstructs the small objects. In Fig. 9(d), DivCon successfully avoids generating the dog outside boxes.
@@
Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy.
4.3.3 Fractiles in Image Generation
During image generation of our DivCon, two fractiles (i.e., $z_{0.3T}$ and $z_{0.7T}$) are used to preserve “good enough” results in the first-round generation. We conduct ablation experiments to study these two fractiles. The second fractile $z_{0.7T}$ marks the end of optimization and is non-adjustable. We observe that transmitting only $z_{0.7T}$ to the next phase results in instability in image generation instability, specifically manifesting as noticeable boundary demarcations of bounding boxes in the images (Fig. 10).
We further compare the performance achieved by our model with different values for the first fractile. Experiment results are presented in Tab. 5. It can be observed that the model suffers a notable performance drop if only the second fractile is employed. With an additional fractile, our DivCon yields a notable improvement. Additionally, the selection of the first fractile’s value exhibits a marginal impact on the efficacy of the model, with the 30% threshold demonstrating optimal performance.
Table 4: Ablation study of our method’s image generation performance on HRS benchmark. "Layout w/DivCon": Layout prediction from LLMs with divide-and-conquer prompting. "Image w/DivCon": Layout-to-Image generation with divide-and-conquer strategy.
Layout w/DivCon Image w/DivCon
Numerical
Precision
Recall
F1
Acc.
Spatial
Acc.
77.43
61.78
68.72
25.60
48.29
78.26
62.61
69.57
29.46
52.74
78.18
62.36
69.38
27.83
53.17
79.13
63.53
70.48
29.62
53.87
Table 5: Ablation study of two fractiles used to preserve results in the first-round generation.
First Fractile Second Fractile
Numerical
Precision
Recall
F1
Acc.
Spatial
Acc.
76.82
60.69
67.80
24.31
48.53
$0.1T$
78.44
61.61
69.01
29.12
52.30
$0.3T$
79.13 63.53 70.48 29.62
53.87
$0.5T$
78.13
63.25
69.90
28.71
51.46
5 Limitations and Discussions
As illustrated in Fig. 11, our method still struggles with generating objects from certain pattern layouts. When the text prompt describes objects overlapping in space, the diffusion model may fail to generate all objects accurately. This limitation stems from the inherent capacity of layout-conditioned image generation models to handle overlapping boxes. Despite these challenges, our approach still demonstrates a reduced frequency of overlapping in layout prediction compared to the base model.
Figure 11: Comparison of transmit one or two latent representations.
6 Conclusion
In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into simple subtasks to achieve higher image generation fidelity. Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, respectively. Then, layout-to-image generation is achieved via an iterative manner to synthesize objects in an easy-to-hard fashion. Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.

### v2

@@ 4.1 Experiment Setup
To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS (Bakr et al. 2023) and NSR-1K (Feng et al. 2023b). The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships. The NSR-1K dataset consists of 762 prompts for numerical relationships and 283 prompts for spatial relationships, all labeled with names, counts, bounding boxes, and ground truth images. Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations. For these benchmarks, Stable Diffusion (Rombach et al. 2022), Attend-and-Excite (Chefer et al. 2023), Divide-and-Bind (Li et al. 2023a), GLIGEN (Li et al. 2023b), Layout-Guidance (Chen, Laina, and Vedaldi 2023b), and Attention-Refocusing (Phung, Ge, and Huang 2024) are included for comparison. In addition, we also include the dataset in (Lian et al. 2023) for evaluation, which includes four tasks: negation, generative numeracy, attribute binding, and spatial reasoning. As our method focuses on numerical and spatial reasoning, we select 100 prompts from these two tasks to construct a benchmark termed LLM-Grounded.
(2) Evaluation Metrics
@@
Quality. To assess the quality of generated images, we use FID as the metric (Heusel et al. 2017). Lower FID represents higher similarity between generated images and ground truths.
@@
Fidelity. To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated. Particularly, YOLOv8 (Jocher, Chaurasia, and Qiu 2023) is employed to obtain the bounding boxes of the objects in the generated images. For counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy (El-Nouby et al. 2019; Fu et al. 2020). For spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills (Cho, Zala, and Bansal 2022) and LayoutGPT (Feng et al. 2023b). Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship. Otherwise, they are denoted as a top-bottom relationship.
Efficiency. To validate the efficiency and scalability of our method, inference runtime is evaluated on an AWS g5.xlarge GPU instance.
@@
Figure 5: Qualitative comparison of numerical & spatial reasoning between Stable Diffusion, Attend-and-excite, Attention -Refocusing and our DivCon.
@@ 4.2 Evaluation Results
The HRS dataset. We present the results on the HRS benchmark in Tab. 1. As we can see, our DivCon outperforms previous T2I models by over 10% in numerical accuracy and about 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance. In addition, DivCon surpasses Attention Refocusing (Phung, Ge, and Huang 2024) on most metrics. Specifically, our layout prediction improves numerical accuracy by 3% and spatial accuracy by 2%, respectively. For image generation, the F1 score and accuracy achieve 2% and 4% gains, while the spatial accuracy is boosted by up to 6%. By dividing the T2I task into simple subtasks, our DivCon can better understand the text prompt to achieve higher fidelity.
@@
The NSR-1K dataset. The results on the NSR-1K benchmark are presented in Tab. 2. It can be observed that DivCon outperforms other T2I models by over 20% in numerical accuracy and about 45% in spatial accuracy. As compared to Attention Refocusing (Phung, Ge, and Huang 2024), our DivCon produces competitive results in terms of layout accuracy. Meanwhile, DivCon achieves approximately 2% gains in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy. Quantitative results in terms of image quality are shown in Tab. 3. As we can see, our DivCon produces comparable FID scores to Stable Diffusion and outperforms attention-refocusing with notable margins. This further demonstrates that our approach is able to produce images with better visual quality while maintaining high prompt fidelity.
Methods FID ($\downarrow$)
Stable Diffusion 2.1 20.77
Attention-Refocusing 21.05
DivCon (Ours) 21.01
Table 3: FID results achieved on NSR-1K.
LLM-Grounded benchmark. Table 4 presents the accuracy achieved by our method and LLM-grounded Diffusion. It is clear that our method achieves comparable performance to LLM-Grounded Diffusion on the numerical reasoning while outperforming it by 3% on the spatial reasoning task. This also validates the superiority of our method.
Methods Numerical Spatial
LLM-Grounded (GPT4) 84% 82%
DivCon (Ours) 84% 87%
Table 4: Comparison with LLM-Grounded
Runtime comparisons are presented in Tab. 5. Among techniques that require multiple forward passes to synthesize different regions separately, our DivCon produces notable efficiency gains. Particularly, LLM-Grounded takes 134.98s to generate an image, while our method achieves a 1.87×1.87$\times$1.87 × speedup. Compared with methods with a single forward pass, our method (1 Iter) produces competitive results with comparable efficiency.
Methods #Forward Pass Runtime
Stable Diffusion 1 15.81 s
GLIGEN 1 33.75 s
Attention-Refocusing 1 35.71 s
DivCon-1 iter (Ours) 1 37.35 s
Mixture-of-Diffusion 3 126.41 s
LLM-Grounded 3 134.98 s
Multi-Diffusion 3 121.68 s
DivCon-2 iter (Ours) 2 72.01 s
Table 5: Comparison of Inference Runtime.
(2) Qualitative Results
@@
Figure 5 illustrates the qualitative comparison of DivCon and baselines. In all cases, our DivCon can accurately predict the object quantities and the spatial relations from the text prompts.
@@
Large Quantity of Objects. Given a prompt containing more than one object category (Fig. 5), Stable Diffusion (Rombach et al. 2022) and Attend-and-Excite (Chefer et al. 2023) often generate an insufficient quantity of objects, while Attention-Refocusing (Phung, Ge, and Huang 2024) sometimes omits one category of objects or generates an excessive number of objects. For instance, the prompt in Fig. 5(b) contains five spoons and four oranges. However, Stable Diffusion and Attend-and-Excite cannot synthesize sufficient spoons or oranges. Meanwhile, Attention-Refocusing generates an excess number of objects, with spoons appearing incomplete due to missing parts. As compared to these methods, our DivCon achieves superior numerical accuracy.
@@
Complicated Spatial Relation. Given a prompt containing complex spatial relationships among more than two object categories, previous T2I models struggle to generate correct spatial relations. For example, Stable Diffusion and Attend-and-Excite miss certain object categories (Fig. 5(d)). Meanwhile, Attention-Refocusing tends to produce distortion or blending artifacts around the boundaries of objects (Fig. 5(c)(d)) or merges the color of the banana with the person (Fig. 5(d)). In contrast, our DivCon produces accurate object categories and spatial relationships, exhibiting superior performance in these challenging cases.
User Study. We conduct a user study to evaluate the images generated by different methods. To be specific, we recruit 17 people with different backgrounds and asked them to rank the images generated by different methods according to their fidelity to the original prompt. Lower ranks indicates better alignment. A total of 20 examples are randomly selected from the NSR-1K benchmark. As shown in Fig. 6, our method achieves the best performance in terms ofaverage rank, which demonstrates the superior perceptual quality of our results.
Figure 6: User Study: Comparison of image quality.
4.3 Ablation Study
In this sub-section, we conduct ablation experiments to evaluate the effectiveness of our divide-and-conquer approach in both layout prediction and image generation. Results are presented in Tab. 6 and Attention-Refocusing (Phung, Ge, and Huang 2024) is used as the baseline.
Divide-and-Conquer in Layout Prediction. To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step. As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 4% in both numerical and spatial accuracy. We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig. 7. Compared to Attention-Refocusing, layouts generated by our DivCon not only exhibit higher reasoning accuracy (Fig. 7 (a) but also are more organized with less overlap (Fig. 7 (b)).
Layout w/ DivCon Image w/ DivCon Numerical Precision Recall F1 Acc. Spatial Acc.
(1) - - 77.43 61.78 68.72 25.60 48.29
(2) - 78.26 62.61 69.57 29.46 52.74
(3) - 78.18 62.36 69.38 27.83 53.17
(4) 79.13 63.53 70.48 29.62 53.87
Table 6: Ablation study of our DiVCon on HRS benchmark. “Layout w/ DivCon”: Layout prediction with divide-and-conquer strategy. “Image w/ DivCon”: Layout-to-Image generation with divide-and-conquer strategy.
Figure 7: Comparison of the predicted layouts.
@@ 4.3 Ablation Study
Divide-and-Conquer in Image Generation. To study the effectiveness of our divide-and-conquer approach in the image generation stage, we develop a variant to synthesize the image in a single forward pass. It can be observed from Tab. 6 that the divide-and-conquer approach facilitates our method to notable performance gains in terms of all metrics. We further compare the results produced by our method using prompts containing multiple objects in Fig. 8. As we can see, the divide-and-conquer approach facilitates our method to better reconstruct the objects. For example, DivCon accurately reconstructs the small objects (Fig. 8(a)) and avoids generating the dog outside boxes (Fig. 8(b)).
@@
Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy (Tab. 6).
4.4 Limitations and Discussions
As illustrated in Fig. 9, when the text prompt describes objects with extreme overlap, the diffusion model may fail to generate all objects accurately. This limitation stems from the inherent capacity of layout-conditioned image generation models to handle overlapping boxes. Despite these challenges, our approach still demonstrates an improved quality in layout prediction as compared to the base model.
Figure 9: An illustration of failure case.
Conclusion
In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into multiple subtasks to achieve higher image generation quality. Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, respectively. Then, layout-to-image generation is achieved via an progressive manner to synthesize objects in an easy-to-hard fashion. Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.